# End-to-End Conformal Calibration for Optimization Under Uncertainty

### v1

@@ 3.1 Conformal uncertainty set calibration
[…]
The split conformal prediction procedure detailed in Lemma 1 gives an approach to selecting $q\in\mathbb{R}$ that ensures the uncertainty set $\Omega_{\theta}(x)$ satisfies $(1-\alpha)$ coverage. Note, however, that this process applies to a given nonconformity score function $s_{\theta}$, and does not address the question of training the uncertainty set $\Omega_{\theta}(x)$ (and hence score function $s_{\theta}$) to ensure optimal task performance while maintaining coverage. We will propose to apply a separate differentiable conformal prediction procedure during training to address this challenge; we discuss this approach in Section 3.3. First, we will describe in the next section two parametric forms of uncertainty set that are commonly used in the literature.
3.2 Representations of the uncertainty set
In this section, we briefly describe two standard uncertainty set representations, box and ellipsoidal uncertainty, which yield tractable forms of the robust optimization problem (1). In both cases, the uncertainty set has the form $\Omega_{\theta}(x)=\left\{\hat{y}\in\mathbb{R}^{n}\mid s_{\theta}(x,\hat{y})\leq q\right\}$ for some function $s_{\theta}:\mathbb{R}^{m}\times\mathbb{R}^{n}\to\mathbb{R}$, and thus matches the form (3) in Section 3.1 that enables conformal calibration. Then, in Section 3.3, we discuss how to train predictive models to generate these uncertainty sets end-to-end within a robust optimization framework.
Box uncertainty sets.
One of the simplest representations of uncertainty is box uncertainty where $\Omega(x)=[\underline{y}(x),\overline{y}(x)]$ is an $n$-dimensional box whose lower and upper bounds depend on $x$. The particular parametrization we use takes inspiration from quantile regression in the univariate setting. Let $h_{\theta}:\mathbb{R}^{m}\to\mathbb{R}^{n}\times\mathbb{R}^{n}$ be a neural network that estimates lower and upper bounds: $h_{\theta}(x)=(h_{\theta}^{\text{lo}}(x),h_{\theta}^{\text{hi}}(x))$. To represent a box uncertainty set in the form (3), we define a nonconformity score function as a generalization of that from scalar conformalized quantile regression [37]:
$s_{\theta}(x,y)=\max\cup_{i=1}^{n}\left\{h_{\theta}^{\text{lo}}(x)_{i}-y_{i},\ y_{i}-h_{\theta}^{\text{hi}}(x)_{i}\right\}.$
Then, the uncertainty set (3) becomes
$\Omega_{\theta}(x)=\{\hat{y}\in\mathbb{R}^{n}\mid s_{\theta}(x,\hat{y})\leq q\}\\ =\left[h_{\theta}^{\text{lo}}(x)-q\mathbf{1},\ h_{\theta}^{\text{hi}}(x)+q\mathbf{1}\right]=:\Big{[}\underline{y}(x),\ \overline{y}(x)\Big{]},$
which is calibrated when $q$ is chosen by the split conformal procedure detailed in Section 3.1.
Given a box uncertainty set $\Omega_{\theta}(x)$, it is possible to take the dual of the inner maximization problem (see Section B.1) to transform the robust optimization problem (1) into an equivalent tractable form:
$z_{\theta}^{\star}(x)=\operatorname*{arg\,min}_{z\in\mathbb{R}^{p}}\min_{\nu\in\mathbb{R}^{n}}$ $\left(\overline{y}(x)-\underline{y}(x)\right)^{\top}\nu+\underline{y}(x)^{\top}Fz+\tilde{f}(x,z)$ (4)
s.t. $\nu\geq\mathbf{0},\quad\nu-Fz\geq\mathbf{0},\quad g(x,z)\leq 0.$
This problem is convex, and hence tractable, so long as $\tilde{f}(x,z)$ and $g(x,z)$ are convex in $z$.
Ellipsoidal uncertainty sets.
Another common form of uncertainty set is that of ellipsoidal uncertainty. Suppose a neural network model outputs predicted mean and covariance parameters $h_{\theta}(x)=(\mu_{\theta}(x),\Sigma_{\theta}(x))$ of a multivariate normal distribution, so that we have a predicted conditional density $\hat{\mathcal{P}}(y\mid x;\theta)=\operatorname{\mathcal{N}}(y\mid\mu_{\theta}(x),\Sigma_{\theta}(x))$, where $\operatorname{\mathcal{N}}(\cdot\mid\mu,\Sigma)$ is the multivariate normal density function. In this case, we define the nonconformity score function based on the squared Mahalanobis distance [22, 41]
which yields uncertainty sets (3) that are ellipsoidal:
$\Omega_{\theta}(x)=\{\hat{y}\in\mathbb{R}^{n}\mid s_{\theta}(x,\hat{y})\leq q\}=\{\hat{y}\in\mathbb{R}^{n}\mid(\hat{y}-\mu_{\theta}(x))^{\top}(\Sigma_{\theta}(x))^{-1}(\hat{y}-\mu_{\theta}(x))\leq q\},$
which is calibrated when $q$ is chosen by the split conformal procedure detailed in Section 3.1. By taking the dual of the inner maximization problem and invoking strong duality (see Section B.2), it is possible to transform the robust optimization problem (1) into the equivalent tractable form:
$z_{\theta}^{\star}(x)=\operatorname*{arg\,min}_{z\in\mathbb{R}^{p}}$ $\sqrt{q}\cdot\|(\Sigma_{\theta}(x))^{1/2}Fz\|_{2}+\mu_{\theta}(x)^{\top}Fz+\tilde{f}(x,z)$ (5)
s.t. $g(x,z)\leq 0,$
which is convex, and hence tractable, so long as $\tilde{f}(x,z)$ and $g(x,z)$ are convex in $z$.
3.3 End-to-end training and calibration
Thus far, we have discussed how to calibrate an uncertainty set $\Omega_{\theta}(x)$ of the form (3) to ensure coverage, and we described two choices of score function $s_{\theta}$ parametrizing common box and ellipsoidal uncertainty sets. However, to ensure that the uncertainty sets $\Omega_{\theta}(x)$ both guarantee coverage and ensure optimal downstream task performance, it is necessary to design an end-to-end training methodology that can incorporate both of these desiderata in a fully differentiable manner. We propose such a methodology in Algorithm 1.
Our end-to-end training approach uses minibatch gradient descent to minimize the empirical task loss $\ell(\theta)=\frac{1}{N}\ell_{i}(\theta)$ where $\ell_{i}(\theta)=f(x_{i},y_{i},z_{\theta}^{\star}(x_{i}))$. This requires differentiating through both the robust optimization problem as well as the conformal prediction step. The gradient of the task loss on a single instance is $\frac{\mathrm{d}\ell_{i}}{\mathrm{d}\theta}=\frac{\partial f}{\partial z}|_{(x_{i},y_{i},z_{\theta}^{\star}(x_{i}))}\frac{\partial z_{\theta}^{\star}}{\partial\theta}|_{x_{i}}$, where $\frac{\partial z_{\theta}^{\star}}{\partial\theta}|_{x_{i}}$ can be computed by differentiating through the Karush–Kuhn–Tucker (KKT) conditions of the convex reformulation of the optimization problem (1) (i.e., the problems (4), (5)) following the approach of Amos and Kolter [3], under mild assumptions on the differentiability of $f$ and $g$. Note that the gradient of any convex optimization problem can be computed with respect to its parameters as such [1, Appendix B].
To include calibration during training, we assume that for every $(x,y)$ in our training set, $s_{\theta}(x,y)$ is differentiable w.r.t. $\theta$ almost everywhere; this assumption holds for common nonconformity score functions, including those used in this paper. We then adopt the conformal training approach [40] in which a separate $q$ is chosen in each minibatch, as shown in Algorithm 1. The chosen $q$ depends on $\theta$ (through $s_{\theta}$), and $z_{\theta}^{\star}(x_{i})$ depends on the chosen $q$. Therefore $\frac{\partial z_{\theta}^{\star}}{\partial\theta}$ involves calculating $\frac{\partial z_{\theta}^{\star}}{\partial q}\frac{\partial q}{\partial\theta}$, where $\frac{\partial q}{\partial\theta}$ requires differentiating through the empirical quantile function. Whereas Stutz et al. [40] uses a smoothed approximate quantile function for calculating $q$, we find the smoothing unnecessary, as the gradient of the empirical quantile function is unique and well-defined almost everywhere. Importantly, our exact gradient is both more computationally efficient and simpler to implement than the smoothed approximate quantile approach. We describe our approach in greater detail in Appendix C.
After training has concluded and we have performed the final conformal calibration step, the resulting model enjoys the following theoretical guarantee on performance (cf. [41, Proposition 1]).
Proposition 1.
After following our training and calibration procedure, we achieve with probability $1-\alpha$ (with respect to $x,y$, and the calibration set $D_{\text{cal}}$) the following upper bound on task loss:
$f(x,y,z_{\theta}^{\star}(x))\leq\min_{z\in\mathbb{R}^{p}}\max_{\hat{y}\in\Omega_{\theta}(x)}f(x,\hat{y},z)\qquad\text{s.t.}\qquad g(x,\hat{y},z)\leq 0.$
This result is an immediate consequence of the split conformal coverage guarantee from Lemma 1, which ensures that for the true $(x,y)\sim\mathcal{P}$, $\mathbb{P}_{x,y,D_{\text{cal}}}(y\in\Omega_{\theta}(x))\geq 1-\alpha$. The realized task loss will therefore, with probability $1-\alpha$, improve on the optimal value of the robust problem (1). Moreover, this guarantee holds despite the fact that the distribution $\mathcal{P}(y\mid x)$ is unknown, and the dependence of the learned uncertainty set $\Omega_{\theta}(x)$ on $x$ allows for taking advantage of heteroskedasticity that, together with our end-to-end training framework, yield substantial improvements in average-case performance in conjunction with our the robust guarantee offered by Proposition 1.
4 Representing General Convex Uncertainty Sets via PICNNs
In the previous section, we discussed how to calibrate a generic family of uncertainty sets $\Omega_{\theta}(x)$ expressed as the sublevel sets of a nonconformity score function as in (3); we then described an end-to-end training methodology for learning such calibrated uncertainty sets to optimize downstream task loss. However, both box and ellipsoidal uncertainty sets are highly restrictive representations of uncertainty whose particular shapes may yield suboptimal task loss performance. If we could instead let $\Omega_{\theta}(x)$ range over the set of all convex uncertainty sets, this more expressive class would enable obtaining better task loss.
To this end, we propose to directly learn a partially-convex nonconformity score function $s_{\theta}:\mathbb{R}^{m}\times\mathbb{R}^{n}\to\mathbb{R}$ that is convex only in the second input vector. Fixing $x$, any $q$-sublevel set $\{\hat{y}\in\mathbb{R}^{n}\mid s_{\theta}(x,\hat{y})\leq q\}$ of $s_{\theta}$ is a convex set, and likewise every family of convex sets can be expressed as the $q$-sublevel sets of some partially-convex function. To implement this approach, we are faced with two questions:
  1. 1.
    How should we parametrize the score function $s_{\theta}$ to allow $\Omega_{\theta}(x)$ to approximate arbitrary convex uncertainty sets?
  2. 2.
    Does the chosen uncertainty set representation yield a tractable convex reformulation of the CRO problem (1)?
A natural answer to the first question is to use a partially input-convex neural network (PICNN) [4] as $s_{\theta}$; these can efficiently approximate any partially-convex function [11]. We consider a PICNN defined as ${s_{\theta}(x,y)=W_{L}\sigma_{L}+V_{L}y+b_{L}}$, where
$\sigma_{0}$ $=\mathbf{0},\qquad u_{0}=x$ $W_{l}$ $=\bar{W}_{l}\operatorname{diag}([\hat{W}_{l}u_{l}+w_{l}]_{+})$ (6a)
$u_{l+1}$ $=\operatorname{ReLU}\left(R_{l}u_{l}+r_{l}\right)$ $V_{l}$ $=\bar{V}_{l}\operatorname{diag}(\hat{V}_{l}u_{l}+v_{l})$ (6b)
$\sigma_{l+1}$ $=\operatorname{ReLU}\left(W_{l}\sigma_{l}+V_{l}y+b_{l}\right)$ $b_{l}$ $=\bar{B}_{l}u_{l}+\bar{b}_{l}.$ (6c)
The weights of the neural network are $\theta=(R_{l},r_{l},\bar{W}_{l},\hat{W}_{l},w_{l},\bar{V}_{l},\hat{V}_{l},v_{l},\bar{B}_{l},\bar{b}_{l})_{l=0}^{L}$, and the matrices $\bar{W}_{l}$ are constrained to be entrywise nonnegative to ensure convexity of $s_{\theta}$ with respect to $y$. For ease of notation, we assume all hidden layers $\sigma_{1},\dotsc,\sigma_{L}$ have the same dimension $d$.
We are now faced with the second question above: does the PICNN representation of the uncertainty set $\Omega_{\theta}(x)$ allow for a tractable reformulation of the CRO problem (1)? Fortunately, as we show in the following theorem, the answer is yes.
Theorem 1.
Let $\Omega_{\theta}(x)=\{\hat{y}\in\mathbb{R}^{n}\mid s_{\theta}(x,\hat{y})\leq q\}$, where $s_{\theta}$ is a PICNN as defined in (6). Then the CRO problem (1) with uncertainty set $\Omega_{\theta}(x)$ is equivalent to the following tractable minimization problem:
s.t. $A(\theta)^{\top}\nu=\begin{bmatrix}Fz\\ \mathbf{0}\end{bmatrix},\quad\nu\geq\mathbf{0},\quad g(x,z)\leq 0$
where $A(\theta)\in\mathbb{R}^{(2Ld+1)\times(n+Ld)}$ and $b(\theta,q)\in\mathbb{R}^{2Ld+1}$ are constructed from the weights $\theta$ of the PICNN (6), and $b$ also depends on $q$. Note that if $\tilde{f}(x,z)$ and $g(x,z)$ are convex in $z$, then (7) is convex, and hence tractable.
We prove Theorem 1 in Section B.3; the main idea is that when $\Omega_{\theta}(x)$ is a sublevel set of a PICNN, we can equivalently reformulate the inner maximization problem in (1) as a linear program, which allows taking the dual of the inner problem to yield a tractable minimization problem.
Since the PICNN uncertainty sets are of the same form as (3) and they yield the tractable convex reformulation (7) of the CRO problem (1), we can apply the same split conformal procedure detailed in Section 3.1 to choose $q\in\mathbb{R}$ and obtain coverage guarantees on $\Omega_{\theta}(x)$, and we can employ the same end-to-end training methodology proposed in Section 3.3 to train calibrated uncertainties end-to-end using the downstream task loss. Note that in some cases during training, the inner maximization problem of (1) with PICNN-parametrized uncertainty set may be unbounded (if $\Omega_{\theta}(x)$ is not compact) or infeasible (if too small a $q$ was chosen by the split conformal procedure and $\Omega_{\theta}(x)$ is empty). This will lead, respectively, to an infeasible or unbounded equivalent problem (7). We can avoid this concern by slightly adjusting the PICNN architecture to ensure its sublevel sets are compact and by suitably increasing $q$ when needed to ensure $\Omega_{\theta}(x)$ is never empty. Such modifications do not change the general form of the problem (7) and preserve the marginal coverage guarantee for the uncertainty set $\Omega(x)$; see Appendix B.4 for details.
5 Experiments
5.1 Problem descriptions
We consider two applications in our experiments: price forecasting for battery storage operation, and a portfolio optimization task.
Price forecasting for battery storage.
This problem comes from Donti et al. [15], where a grid-scale battery operator predicts electricity prices $y\in\mathbb{R}^{T}$ over a $T$-step horizon and uses the predicted prices to decide on a battery charge/discharge schedule for price arbitrage. The input features $x$ include the past day’s prices and temperature, the next day’s energy load forecast and temperature forecast, binary indicators of weekends or holidays, and yearly sinusoidal features. The operator decides how much to charge ($z^{\mathrm{in}}\in\mathbb{R}^{T}$) or discharge ($z^{\mathrm{out}}\in\mathbb{R}^{T}$) the battery, which changes the battery’s state of charge ($z^{\mathrm{state}}\in\mathbb{R}^{T}$). The battery has capacity $B$, charging efficiency $\gamma$, and maximum charging/discharging rates $c^{\mathrm{in}}$ and $c^{\mathrm{out}}$. The task loss function represents the multiple objectives of maximizing profit, flexibility to participate in other markets by keeping the battery near half its capacity (with weight $\lambda$), and battery health by discouraging rapid charging/discharging (with weight $\epsilon$):
The constraints are given by
$z^{\mathrm{state}}_{0}=B/2,\qquad z^{\mathrm{state}}_{t}=z^{\mathrm{state}}_{t-1}-z^{\mathrm{out}}_{t}+\gamma z^{\mathrm{in}}_{t}\qquad\forall t=1,\dotsc,T$
$0\leq z^{\mathrm{in}}\leq c^{\mathrm{in}},\qquad 0\leq z^{\mathrm{out}}\leq c^{\mathrm{out}},\qquad 0\leq z^{\mathrm{state}}\leq B.$
Following Donti et al. [15], we set $T=24$ hours, $B=1$, $\gamma=0.9$, $c^{\mathrm{in}}=0.5$, $c^{\mathrm{out}}=0.2$, $\lambda=0.1$, and $\epsilon=0.05$.
Portfolio optimization
[…]
Following the experimental setup in Chenreddy and Delage [12], we consider a synthetic dataset where $(x,y)\in\mathbb{R}^{2\times 2}$ are drawn from a mixture of three 4-D multivariate normal distributions. Due to space constraints, we provide more data details in Section D.2 and report experimental results in Section D.3.
5.2 Baseline methods
We implemented several “estimate-then-optimize” (ETO) baselines, listed below, to compare against our end-to-end (E2E) method. These two-stage ETO baselines are trained using task-agnostic losses such as pinball loss or negative log-likelihood, with conformal calibration applied after training to satisfy marginal coverage and ensure fair comparison against our E2E method.
  • ETO denotes models that have identical neural network architectures to our E2E models, differing only in the loss function during training. The box uncertainty ETO model is trained with pinball loss to predict the $\alpha/2$ and $21-\alpha/$ 2 quantiles. The ellipsoidal uncertainty ETO model predicts a mean vector and covariance matrix and is trained with a multivariate normal negative log-likelihood loss. For the PICNN ETO model, we give probabilistic interpretation to a PICNN model $s_{\theta}$ via the energy-based model $\hat{\mathcal{P}}(y\mid x)\propto\exp(-s_{\theta}(x,y))$ and train it using the negative log-likelihood (NLL) loss function $\operatorname{NLL}(s_{\theta}(x,y))=\ln s_{\theta}(x,y)+\ln Z_{\theta}(x),$ where $Z_{\theta}(x)=\int_{y\in\mathbb{R}^{n}}\exp(-s_{\theta}(x,y))\,\mathop{}\!\mathrm{d}y$, following the approach of Lin and Ba [26]. More details of the ETO models are given in Appendix D.
  • ETO-SLL denotes our implementation of the box and ellipsoid uncertainty ETO methods from Sun et al. [41]. Unlike our ETO models, ETO-SLL first trains a point estimate model (without uncertainty) with mean-squared error loss. Then, ETO-SLL box and ellipsoidal uncertainty sets are derived from training a separate quantile regressor using pinball loss to predict the $(1-\alpha)$-quantiles of absolute residuals or $\ell_{2}$-norm of residuals of the point estimate. Unlike ETO which can learn ellipsoidal uncertainty sets with different covariance matrices for each input $x$, the ETO-SLL ellipsoidal uncertainty sets all share the same covariance matrix (scaled differently for each $x$).
  • ETO-JC denotes our implementation of the ETO method by Johnstone and Cox [22], which only applies to ellipsoid uncertainty. Like ETO-SLL, ETO-JC also first trains a point estimate model (without uncertainty) with mean-squared error loss. ETO-JC uses the same covariance matrix (with the same scale) for each input $x$.
5.3 Experimental results: Battery storage problem
Figure 2: Task loss performance (mean $\pm$1 stddev across 10 runs) for the battery storage problem with no distribution shift. Lower values are better.
Figure 3: Coverage (mean $\pm$1 stddev across 10 runs) for the battery storage problem with no distribution shift, with the dotted black line indicating the target coverage level $1-\alpha$.
[…]
5.4 Performance under distribution shift
Figure 4: Task loss performance (mean $\pm$1 stddev across 10 runs) for the battery storage problem with distribution shift. Lower values are better.
Figure 5: Coverage (mean $\pm$1 stddev across 10 runs) for the battery storage problem with distribution shift, with the dotted black line indicating the target coverage level $1-\alpha$.
The aforementioned results were produced without distribution shift, where our training and test sets were sampled uniformly at random, thus ensuring exchangeability and guaranteeing marginal coverage. In this section, we evaluate our method on the more realistic setting with distribution shift by splitting our data temporally; our models are trained on the first 80% of days and evaluated on the last 20% of days. Figures 5 and 5 mirror Figures 3 and 5, except that there is now distribution shift. We again find that our E2E approach consistently yields improved performance over all ETO baselines, for all three uncertainty set parameterizations, and for all tested uncertainty levels $\alpha$. Moreover, both the ellipsoidal and PICNN uncertainty representations, when trained end-to-end, improve on the performance offered by box uncertainty, with PICNN and ellipsoidal uncertainty performing comparably. In terms of coverage (Figure 5), we find that under distribution shift, the models do not provide the same level of coverage guaranteed in the i.i.d. case, and the ellipsoidal and PICNN models tend to provide worse coverage than the box uncertainty. This is not too surprising, however, since these two models offer greater representational power, and thus might be fitting too closely to the pre-shift distribution, which impacts robustness under distribution shift. Devising methods to anticipate distribution shift when training these more expressive models, and in particular the PICNN-based uncertainty, remains an interesting avenue for future work.
6 Conclusion
In this work, we develop the first end-to-end methodology for training predictive models with uncertainty estimates (with calibration enforced differentiably throughout training) that are utilized in downstream conditional robust optimization problems. We propose an approach utilizing partially input-convex neural networks (PICNNs) to represent general convex uncertainty regions, and we perform extensive experiments on a battery storage application and a portfolio optimization task, which validate the utility of our end-to-end approach and the expressiveness of the PICNN representation.
Limitations
Our framework requires various assumptions on the functions $f$ and $g$ in the conditional robust optimization problem (1) in order to ensure tractability, such as convexity and bilinearity in $(y,z)$. In addition, while the PICNN uncertainty representation can represent general convex uncertainties, it cannot handle more general nonconvex uncertainty regions. Removing these assumptions and approaching more general stochastic and robust optimization problems using, e.g., policy gradient methods would be an interesting direction for future work.
Broader Impacts
Our work, which introduces a methodology to improve the performance and robustness for a class of robust optimization problems, has the potential for positive societal impact via deployment to problems such as battery storage operation, where better decisions in the face of uncertainty can help with grid-scale decarbonization efforts. Due to the primarily methodological contributions of this work, we do not anticipate negative societal impacts or the need for safeguards.
Acknowledgments
We thank Priya Donti for helpful discussions. The authors acknowledge support from an NSF Graduate Research Fellowship (DGE-2139433); NSF Grants CNS-2146814, CPS-2136197, CNS-2106403, and NGSDI-2105648; Amazon AWS; and the Resnick Sustainability Institute.
Appendix A Appendix: Related Work
Task-based learning.
The notion of “task-based” end-to-end model learning was introduced by Donti et al. [15], which proposed to train machine learning models end-to-end in a manner capturing a downstream stochastic optimization task. To achieve this, the authors backpropagate gradients through a stochastic optimization problem, which is made possible for various types of convex optimization problems via the implicit function theory [15, 3, 1]. However, Donti et al. [15] does not train the model to estimate uncertainty and thereby does not provide any explicit guarantees on robustness on their decisions to uncertainty. Our framework improves upon this baseline by yielding calibrated uncertainty sets which can then be used to obtain robust decisions.

### v2

@@ A.1 Experimental results: Battery storage
We note that in theory, E2E times should always be higher than the optimization time. The optimization time accounts for the time required to compute $z_{\theta}^{\star}(x)$, and E2E training additional accounts for the time required to compute the gradient $\frac{\partial}{\partial\theta}z_{\theta}^{\star}(x)$. However, in practice, the optimization time depends heavily on the choice of numerical solver. In our implementation, we use the default cvxpy solver (Clarabel) for the optimization step in ETO, whereas we use the default cvxpylayers solver (SCS) during E2E training. In the case of the Ellipse and PICNN uncertainty sets, the SCS solver is generally faster than Clarabel; hence, Table˜A1 counterintuitively reports lower E2E training times for Ellipse and PICNN than optimization times.
We further note that the optimization and E2E training times reflect our particular implementation, but significantly faster times are likely possible. In particular, we solve all of the optimization problems (in both the ETO optimization step and during E2E training) using CPU-based solvers; however, recently developed GPU-accelerated solvers may be able to reduce the optimization time required by orders of magnitude, especially when solving a batch of problems with the same structure (blin_accelerate_2024).
no distribution shift
with distribution shift
Figure A1: Similar to Figure˜3, except that the value plotted is the VaR[$^{1-\alpha}[\text{task loss}$] (mean $\pm$1 stddev across 10 runs) on the test set for the battery storage problem with no distribution shift (top) and with distribution shift (bottom). Lower values are better.
no distribution shift
with distribution shift
Figure A2: Similar to Figure˜3, except that the value plotted is the CVaR[$^{1-\alpha}[\text{task loss}$] (mean $\pm$1 stddev across 10 runs) on the test set for the battery storage problem with no distribution shift (top) and with distribution shift (bottom). Lower values are better.
Table A1: Time per epoch of pretraining, optimization, and end-to-end training, measured in seconds (wall-clock time). Values are mean $\pm$1 stddev across 10 epochs. Lower values are better.
pretrain optimize (train+val) E2E training
Box 0.03 $\pm$ 0.01 4.29 $\pm$ 0.07 6.34 $\pm$ 0.32
Ellipse 0.03 $\pm$ 0.01 7.79 $\pm$ 0.06 7.12 $\pm$ 1.28
PICNN 4.20 $\pm$ 0.20 50.49 $\pm$ 0.19 33.24 $\pm$ 0.25
A.2 Experimental results: Portfolio optimization
Tables˜A2 and A3 show the task loss and coverage results for the portfolio optimization problem. We again find that our E2E approach generally improves upon the ETO baselines at all uncertainty levels $\alpha$, with the exception of box uncertainty, where all the methods achieve similar performance. Our PICNN-based uncertainty representation, when learned end-to-end, performs better than box uncertainty and comparably with ellipse uncertainty. The similarity in performance between E2E ellipsoidal uncertainty and E2E PICNN uncertainty is likely due to the underlying aleatoric uncertainty (i.e., the uncertainty in $\mathcal{P}(y\mid x)$) generally taking an ellipsoidal shape—the conditional distribution $\mathcal{P}(y\mid x)$ is a Gaussian mixture model, and it tends to have a dominant mode (e.g., see Figure˜A3). In terms of coverage, we find that all the models and training methodologies obtain coverage very close to the target level, confirming that the improvements in task loss performance from our E2E approach do not come at the cost of worse coverage.
Because the conditional distribution $\mathcal{P}(y\mid x)$ for the portfolio optimization problem is 2-dimensional, we can visualize the conditional distribution as well as the uncertainty sets estimated by our models. Figure˜A3 plots the conditional density for input $x=\begin{bmatrix}-1.167&0.024\end{bmatrix}^{\top}$, along with the $\alpha=0.1$ uncertainty sets $\Omega_{\theta}(x)$ and the resulting decision vectors $z_{\theta}^{\star}(x)$ for each uncertainty set parametrization. Uncertainty sets and decision vectors from both ETO and E2E models are shown in different colors. The key takeaway from this figure is that smaller uncertainty sets (which is what ETO training tends to produce) do not always result in lower task loss. Furthermore, the more flexible parametrization of the PICNN allows it to learn uncertainty set shapes that may be more amenable to the downstream robust decision task than box or ellipsoidal uncertainty, even if the resulting uncertainty set has a larger or odder shape.
Table A2: Task loss performance (mean $\pm$ 1 stddev across 10 runs) for the portfolio optimization problem. Lower values are better, and the best performance for each uncertainty-level $\alpha$ is highlighted. The results show that our E2E methods consistently outperform the ETO baselines.
uncertainty level $\alpha$
0.01 0.05 0.1 0.2
ETO Box -1.16 $\pm$ 0.42 -1.37 $\pm$ 0.12 -1.39 $\pm$ 0.13 -1.41 $\pm$ 0.12
ETO Ellipse -1.09 $\pm$ 0.12 -1.24 $\pm$ 0.11 -1.29 $\pm$ 0.10 -1.33 $\pm$ 0.10
ETO PICNN -0.95 $\pm$ 0.24 -1.11 $\pm$ 0.24 -1.20 $\pm$ 0.22 -1.31 $\pm$ 0.16
ETO-SLL Box -1.41 $\pm$ 0.13 -1.42 $\pm$ 0.12 -1.42 $\pm$ 0.12 -1.44 $\pm$ 0.11
ETO-SLL Ellipse -1.12 $\pm$ 0.22 -1.37 $\pm$ 0.12 -1.40 $\pm$ 0.12 -1.43 $\pm$ 0.12
ETO-JC Ellipse -1.16 $\pm$ 0.17 -1.40 $\pm$ 0.11 -1.42 $\pm$ 0.11 -1.44 $\pm$ 0.11
E2E Box -1.21 $\pm$ 0.44 -1.40 $\pm$ 0.14 -1.43 $\pm$ 0.11 -1.43 $\pm$ 0.10
E2E Ellipse -1.48 $\pm$ 0.12 -1.47 $\pm$ 0.11 -1.48 $\pm$ 0.11 -1.47 $\pm$ 0.11
E2E PICNN -1.45 $\pm$ 0.14 -1.48 $\pm$ 0.10 -1.48 $\pm$ 0.10 -1.47 $\pm$ 0.11
Table A3: Coverage (mean $\pm$1 stddev across 10 runs) for the portfolio optimization problem. Our E2E models achieve similar coverage to the ETO baselines, confirming that the lower task loss of our E2E models does not come at the expense of worse coverage.
uncertainty level $\alpha$
0.01 0.05 0.1 0.2
ETO Box .984 $\pm$ .007 .947 $\pm$ .017 .902 $\pm$ .017 .786 $\pm$ .020
ETO Ellipse .988 $\pm$ .004 .944 $\pm$ .020 .894 $\pm$ .022 .794 $\pm$ .027
ETO PICNN .989 $\pm$ .006 .949 $\pm$ .014 .901 $\pm$ .019 .801 $\pm$ .034
ETO-SLL Box .985 $\pm$ .012 .945 $\pm$ .021 .885 $\pm$ .030 .796 $\pm$ .029
ETO-SLL Ellipse .989 $\pm$ .011 .945 $\pm$ .024 .885 $\pm$ .039 .795 $\pm$ .030
ETO-JC Ellipse .991 $\pm$ .006 .953 $\pm$ .017 .902 $\pm$ .026 .796 $\pm$ .026
E2E Box .989 $\pm$ .006 .949 $\pm$ .012 .903 $\pm$ .016 .785 $\pm$ .019
E2E Ellipse .992 $\pm$ .006 .954 $\pm$ .010 .903 $\pm$ .022 .798 $\pm$ .022
E2E PICNN .993 $\pm$ .002 .953 $\pm$ .010 .912 $\pm$ .017 .798 $\pm$ .024
Figure A3: This figure plots the density of the conditional distribution $\mathcal{P}(y\mid x)$ for $x=\begin{bmatrix}-1.167&0.024\end{bmatrix}^{\top}$ from the portfolio optimization problem, with darker colors indicating higher density. Also plotted are the $\alpha=0.1$ uncertainty sets $\Omega_{\theta}(x)$ (dashed lines) and the resulting decision vectors $z_{\theta}^{\star}(x)$ (arrows) for each uncertainty set parametrization. Results for ETO models are shown in blue, whereas results for E2E are shown in orange. The “true” $y$ sampled from $\mathcal{P}(y\mid x)$ is drawn in green, and the task loss for this example is computed using this $y$. The decision vectors have been artificially scaled larger to be easier to see.
Appendix B Appendix: Related Work
“Task-based” or “decision-focused” learning.
The notion of “task-based” end-to-end model learning was introduced by donti_task-based_2017, which proposed to train machine learning models in an end-to-end fashion to minimize a downstream stochastic optimization objective. To achieve this, the authors backpropagate gradients through a stochastic optimization problem, which is made possible for various types of convex optimization problems via the implicit function theorem (donti_task-based_2017; amos_optnet_2017; agrawal_differentiable_2019). However, donti_task-based_2017 does not train the model to estimate uncertainty and thereby does not provide any explicit guarantees on robustness on their decisions to uncertainty. Our framework improves upon this baseline by yielding calibrated uncertainty sets which can then be used to obtain robust decisions.
Another line of related works relies on surrogate loss functions to approximate the gradient of the downstream task loss, typically in settings where the exact gradient does not exist or is otherwise hard to compute. For example, the “Smart Predict, then Optimize” approach (elmachtoub_smart_2022) specifically considers differentiating the solution of linear programs, whereas wilder_melding_2019 differentiates the solution of discrete (combinatorial) optimization problems. As with donti_task-based_2017, though, these approaches focus only on average task loss without any estimation of model uncertainty, thereby lacking explicit guarantees on the robustness of the resulting decisions. A recent survey by mandi_decision-focused_2024 provides a comprehensive review of existing decision-focused learning approaches.